# Admissibility of Stein Shrinkage for Batch Normalization in the Presence of Adversarial Attacks

### v1

@@ Abstract
Batch normalization (BN) is a ubiquitous operation in deep neural networks used primarily to achieve stability and regularization during network training. BN involves feature map centering and scaling using sample means and variances, respectively. Since these statistics are being estimated across the feature maps within a batch, this problem is ideally suited for the application of Stein’s shrinkage estimation, which leads to a better, in the mean-squared-error sense, estimate of the mean and variance of the batch. In this paper, we prove that the Stein shrinkage estimator for the mean and variance dominates over the sample mean and variance estimators in the presence of adversarial attacks when modeling these attacks using sub-Gaussian distributions. This facilitates and justifies the application of Stein shrinkage to estimate the mean and variance parameters in BN and use it in image classification (segmentation) tasks with and without adversarial attacks. We present SOTA performance results using this Stein corrected batch norm in a standard ResNet architecture applied to the task of image classification using CIFAR-10 data, 3D CNN on PPMI (neuroimaging) data and image segmentation using HRNet on Cityscape data with and without adversarial attacks.
@@ 1 Introduction
Deep Neural Networks are now widely accepted and used as the primary workhorse in many of the complicated function approximation tasks encountered in engineering and sciences. In particular, the field of computer vision has seen a major paradigm shift from traditional innovative feature engineering to data-driven feature learning via the use of deep neural networks. Deep network training is well known to be time intensive and unstable. Batch normalization (BN) is a particularly useful tool in achieving numerical stability and regularization during network training. BN typically involves centering and scaling the feature maps. The centering and scaling operations involve computing the sample means and variances, respectively, from the samples/members in the batch. This task is ideally suited for the application of Stein’s shrinkage method, which is well known in the field of statistics. Before delving into the Stein’s shrinkage method, let us briefly review the widely popular batch normalization in deep learning.
@@
Batch normalization was introduced in the machine learning literature primarily to accelerate training and reduce the so-called internal covariate shift Ioffe-ICML15 . Since its inception, several variants have populated the literature, including, but not limited to, layer normalization in Recurrent Neural Networks (RNNs) ba2016layernormalization and Transformers, spectral normalization in generative adversarial networks (GANs) Miyato-ICLR18 , and many others. Improvements over BN in the form of a stochastic “whitening” transform to decorrelate the features in each layer at each step of the training has also been proposed in Zhang_2021_CVPR and batch normalization preconditioning in Lange-JMLR2022 , wherein the pre-conditioner is constructed using batch statistics. These normalization algorithms do not address the effectiveness of improvements in stability or convergence of training in the presence of adversarial attacks.
In this paper, we address the problem of developing a BN that leads to better prediction accuracy from the deep network used in classification in the presence of adversarial attacks. Adversarial attacks have been modeled via additive bounded perturbations MauryaArxiv22 . Some of the popular adversarial attack methods that are popular and satisfy the bounded perturbation condition are FGSM (fast gradient sign method) szegedy2014 and an iterative variaint of PGD (projected gradient descent). Numerous articles exist in the literature that depict ways to overcome a drop in the accuracy of classifiers in the presence of adversarial attacks. Our goal here is not the development of a statistically robust estimator that is insensitive to adversarial attacks by seeking a robust loss or modifying the network architecture, however, to adapt an existing shrinkage estimator specifically, the well known James-Stein shrinkage estimator and prove that it retains its dominance over the sample mean and variance in the presence of adversarial attacks. Dominance here refers to smaller mean squared error (MSE) – between the estimate and the true value – when compared to the sample mean and variance. This smaller MSE implies a better BN in each layer and as we demonstrate empirically, a better stability and prediction accuracy from the deep network. By making BN stable in the presence of adversarial attacks this work could have potential applications that might have societal impact.
1.1 A Brief on the James-Stein Shrinkage Estimator
@@
In the following we present a brief introduction to the James-Stein (JS) shrinkage estimator and motivate it’s aptness in the context of batch normalization in deep neural networks. This will provide the necessary background and motivation for our work presented in this paper.
Shrinkage estimation is a parameter estimation technique that is primarily used for variance reduction and to improve parameter estimation accuracy by shrinking the parameter towards a fixed target. For instance, shrinking the estimate of the mean of a Gaussian toward zero. It is especially useful in high-dimensional and scarce data settings. The shrinkage, of course, leads to a biased estimate, but with smaller variance, which is sought after in many applications. The James-Stein (JS) shrinkage estimator was proposed in the field of statistics as an estimator of the mean for samples drawn from a Gaussian distribution and has been shown to dominate the maximum likelihood estimator (MLE) in terms of the risk (mean-squared error) James-Stein61 . Dominance here refers to the risk associated with the JS-estimator being at most as much as that of the MLE and that there exists at least one set of parameter values for which it is strictly less. More precisely, if $\mathbf{X}$ is a $p$=dimensional random vector whose mean is to be estimated and it is assumed that $\mathbf{X}\sim N(\mathbf{\mu},\mathbf{I})$, then, $E(||\mathbf{\mu}^{JS}-\mathbf{\mu}||^{2})\leq E(||\mathbf{\mu}^{MLE}-\mathbf{\mu}||^{2})$. It is important to note that the JS-estimator improves the overall risk of estimating the mean (parameter) vector and not the risk of the individual elements of the mean vector.
@@
At the foundation of the JS-shrinkage estimator lies what is known as Stein’s (identity) lemma, which characterizes the normal distribution. Stein’s (identity) lemma is also at the foundation of the Stein’s method which is used for distributional comparisons. We will not discuss this latter well-known method in statistics here and instead refer the interested reader to a recent comprehensive survey Anastasiou23 .
@@
Since its inception, the JS-shrinkage estimator has been generalized and new optimal shrinkage estimators have been developed for application to many types of distributions such as the location family Brown66 , exponential family Hudson78AOS , Gamma distribution DasGupta86AOS , distributions with quadratic varaince function XianchaoXie-Brown-AOS16 and FathiAOS2022 for relaxing the Gaussian assumption in high dimensions. For a comprehensive survey on Stein’s method and shrinkage estimators, we refer the reader to Kubokawa24 . In addition, more recently, Stein’s shrinkage estimate has been generalized to distributions on manifold-valued random variables Yang-Doss-Vemuri22 .
@@ 1.2 Relevance of JS-estimator to BN
Given the above brief introduction to the JS-shrinkage estimator, we are now ready to present it’s relevance to batch normalization in deep neural networks and how it can be applied to improve the stability and accuracy of predictions. A wide range of normalization techniques have been developed to facilitate the training of deep neural networks by normalizing intermediate feature distributions and improving the stochastic optimization dynamics. One of the earliest and most influential approaches is BN, introduced in Ioffe-ICML15 , which normalizes features using statistics computed over the entire mini-batch. This method was originally motivated by the desire to reduce internal covariate shift, though subsequent work has highlighted its strong impact on the optimization landscape conf/nips/SanturkarTIM18 as a key factor behind its effectiveness. In the following, we give a brief and precise description of different types of normalization that are in vogue in literature.
Let $X\in\mathbb{R}^{N\times C\times H\times W}$ denote the feature tensor at a given layer of a deep neural network, where $N$ – batch size, $C$ – number of channels (feature maps), $H,W$ – spatial dimension. Formally, many normalization techniques follow a common template. Each element $x_{i}\in X$ is normalized as: $\hat{x_{i}}=\frac{x_{i}-\mu_{i}}{\sqrt{\sigma^{2}_{i}+\epsilon}}$, where $i=(i_{N},i_{C},i_{H},i_{W})$ is a multi-index ranging over the tensor dimensions and $\mu_{i},\sigma^{2}_{i}$ are the empirical mean and variance computed over a subset $S_{i}\subseteq X:$ $\mu_{i}=\frac{1}{|S_{i}|}\sum_{k\in S_{i}}x_{k},\ \ \sigma_{i}^{2}=\sum_{k\in S_{i}}(x_{k}-\mu_{i})^{2}.$ The choice of $S_{i}$ determines the normalization scheme and the available choices are the following: (i) Batch Normalization (BN)  Ioffe-ICML15 : $S_{i}^{BN}={(n^{\prime},h^{\prime},w^{\prime})|n^{\prime}\in[1,N],h^{\prime}\in[1,H],w^{\prime}\in[1,W]}.$ Statistics are computed over all samples in the batch and all spatial positions, but separately for each channel $i_{c}.$ (ii) Layer Normalization (LN)  ba2016layernormalization : $S_{i}^{LN}={(c^{\prime},h^{\prime},w^{\prime})|c^{\prime}\in[1,C],h^{\prime}\in[1,H],w^{\prime}\in[1,W]}.$ Statistics are computed independently for each sample (fixed $i_{N}$) across all channels and spatial locations. (iii) Instance Normalization (LN)  ulyanov2017instancenormalizationmissingingredient : $S_{i}^{IN}={(h^{\prime},w^{\prime})|h^{\prime}\in[1,H],w^{\prime}\in[1,W]}.$ Normalization is per-sample and per-channel — only over the spatial dimensions. (iv) Group Normalization (GN)  wu2018groupnormalization : $S_{i}^{GN}={(c^{\prime},h^{\prime},w^{\prime})|c^{\prime}\in C_{\text{group}},h^{\prime}\in[1,H],w^{\prime}\in[1,W]}.$ Channels are split into groups; statistics are computed over spatial dimensions and within the channel group. After normalization, a learnable affine transformation is applied to preserve representational flexibility: $y_{i}=\gamma_{i}\hat{x_{i}}+\beta_{i},$ where $\gamma_{i},\beta_{i}$ are trainable parameters indexed by coordinates across which we compute the statistics.
@@
While normalization methods like LayerNorm, InstanceNorm, and GroupNorm have been introduced to mitigate batch-size dependence, BatchNorm remains a central component in many modern architectures and continues to attract interest both in theory and practice. A key observation is that BatchNorm estimates a large number of mean and variance values simultaneously — one pair for each channel. This can be naturally viewed as a multitask estimation problem: we are estimating a vector of means $\mu=(\mu_{1},\dots,\mu_{c})$ and variances $\sigma^{2}=(\sigma^{2}_{1},\dots\sigma^{2}_{c})$ each based on a finite number of observations.
@@
In such settings, James–Stein shrinkage estimators offer a statistically grounded way to improve estimation by borrowing strength across tasks — in this case, across channels. The classical result suggests that when estimating multiple parameters jointly under squared loss, shrinking individual estimates toward their grand mean can reduce overall error. Applying this idea in the context of BatchNorm could lead to more accurate and stable normalization statistics, particularly in small batch or high variance regimes, where the empirical estimates of mean and variance are noisy and benefit from regularization.
@@
Having established adequate motivation for application of Stein’s shrinkage to BN in a deep network, we are now ready to narrate the rest of the story. But before that, a word on what is to follow. Section 2 delves into the presentation of the appropriate shrinkage formulas for the BN. This is followed by the presentation of theoretical results on the dominance of Stein’s shrinkage formulas in the presence of adversarial attacks in Section 3. In Section 4, we present experimental results that compare the performance of the proposed shrinkage to SOTA on several benchmark data sets. Finally in Section 5, we present the discussion and conclusions.
2 Stein’s Shrinkage Applied to Batch Normalization
In this section, we apply the Stein’s shrinkage to the BN and present the formulas for calculation of the batch normalization in a deep neural network. We will limit ourselves primarily to convolutional neural networks (CNNs) but the formulas can be applied to any architecture with some minor changes.
@@ 2 Stein’s Shrinkage Applied to Batch Normalization
Let $X\sim\mathcal{N}_{p}(\theta,I_{p})$, where $\theta\in\mathbb{R}^{p}$ is an unknown parameter vector and $I_{p}$ is the $p$-dimensional identity matrix. The James Stein shrinkage estimator of the unknown parameter $\theta$ is given by, $\hat{\theta}^{\mathrm{JS}}=\left(1-\frac{p-2}{\|X\|^{2}_{2}}\right)X$. It is well known that JS-Estimator has a smaller mean-squared error than the maximum likelihood estimator but is obviously a biased estimator. However, a smaller mean squared error is desirable in many applications. Most recently, JS-estimator for the calculation of the mean in BN was used in khoshsirat2023improvingnormalizationjamessteinestimator . However, the authors in khoshsirat2023improvingnormalizationjamessteinestimator used the same JS-estimator formula to shrink the mean as well as the variance in BN. Directly applying the same shrinkage formula to the variance parameter is, however, theoretically flawed, as the sample variance is not normally distributed but instead follows a scaled chi-squared distribution - a special case of the Gamma distribution. To address this, we adopt a principled alternative.
More specifically, for mean correction, we adopt the James-Stein shrinkage under the assumption that individual elements of the feature tensor in the BN are normally distributed and homoscedastic. The homoscedasicity assumption simplifies our calculations and we will address relaxing this assumption in future. Under this assumption, the sample mean is normally distributed, and the sample variance follows a chi-squared distribution (i.e., a Gamma distribution): $\hat{\sigma}_{i}^{2}\sim\frac{\sigma_{i}^{2}}{n}\cdot\chi^{2}_{n-1}\sim\mathrm{Gamma}\left(\alpha=\frac{n-1}{2},\ \beta_{i}=\frac{2\sigma_{i}^{2}}{n}\right)$, which enables us to apply shrinkage to the scale parameter of the Gamma distribution DasGupta86AOS . Let $X_{i}\sim\mathrm{Gamma}(\alpha,\beta_{i})$ independently for $i=1,\dots,p$, where all shape parameters are equal: $\alpha_{1}=\dots=\alpha_{p}=\alpha$. Then JS-estimators for the scale parameters $\beta=(\beta_{1},\dots,\beta_{p})\in\mathbb{R}_{+}^{p}$ are as follows:
@@
$\hat{\beta}_{i}^{\,\mathrm{JS}}=\frac{X_{i}}{\alpha+1}+\tilde{c}V,\quad\text{where}\quad V=\left(\prod_{j=1}^{p}X_{j}\right)^{1/p},\tilde{c}\in\left[0,\dfrac{2(p-1)}{(\alpha+1)(\alpha p+1)}\right]$ (1)
@@
To incorporate the James-Stein correction into Batch Normalization, we compute the statistics for each channel and apply the shrinkage formulas as outlined below. Given a 4D tensor input $x\in\mathbb{R}^{N\times C\times H\times W},$ we define $\mu_{C}\in\mathbb{R}^{C}$ and $\sigma^{2}_{C}\in\mathbb{R}^{C}_{+}$ as the channel-wise mean and variance computed over all spatial and batch dimensions given by:
@@
We then apply shrinkage to these estimates. The James-Stein corrected mean and variance are:
@@
Equations 2 and 3 respectively are the formulas we advocate for use in BN corrections and will be used in all our experiments. We are now ready to present the theoretical results of dominance of these shrinkage estimators in the presence of adverserial attacks.
3 Dominance of Stein Shrinkage in the Presence of Adverserial Attacks
In this section, we present the main theoretical contribution of our work: the James–Stein estimator continues to exhibit dominance even under adversarial perturbations. We begin by formalizing how adversarial attacks can be modeled probabilistically. Then, we introduce the necessary preliminaries and outline the main ideas behind the proof.
We model adversarial perturbations using a sub-Gaussian distribution, which is well-motivated both practically and theoretically. Adversarial perturbations are typically bounded in magnitude (e.g., under an $l_{p}$-norm constraint) to ensure the model remains learnable — a standard assumption in the literature cohen2019certifiedadversarialrobustnessrandomized ; zhai2019adversariallyrobustgeneralizationjust . Since bounded random variables are inherently sub-Gaussian, this provides a natural modeling choice. Additionally, the sub-Gaussian model enables a tractable analysis via the use of concentration inequalities and allows one to work with a well-defined variance proxy. Following prior workMauryaArxiv22 , we assume that each adversarial perturbation vector has coordinates bounded in $[-\varepsilon,\varepsilon]$ and can be represented as a zero-mean sub-Gaussian vector with variance proxy $2\varepsilon^{2}$: $Y\sim\text{subG}(2\varepsilon^{2}),\quad\text{i.e., for all }v\in\mathbb{R}^{p}\text{ with }\|v\|=1,\ \ \text{and}\ \ \mathbb{E}\left[\exp\left(\lambda\langle v,Y\rangle\right)\right]\leq\exp\left(\frac{2\varepsilon^{2}\lambda^{2}}{2}\right),\quad\forall\lambda\in\mathbb{R}.$ This assumption allows us to model the observed vector as the sum of noise-free data and adversarial noise: $Z=X+Y,\ \ \ X\sim\mathcal{N}_{p}(\theta,\sigma^{2}I).$ For simplicity we assume that $X$ and $Y$ are independent, which is currently a limitation of our model. Other noise models that are data dependent will be considered in our future work.

### v2

@@ Abstract
Batch normalization (BN) is a ubiquitous operation in deep neural networks, primarily used to improve stability and regularization during training. BN centers and scales feature maps using sample means and variances, which are naturally suited for Stein’s shrinkage estimation. Applying such shrinkage yields more accurate mean and variance estimates of the batch in the mean-squared-error sense. In this paper, we prove that the Stein shrinkage estimator of the mean and variance dominates over the sample mean and variance estimators, respectively, in the presence of adversarial attacks modeled using sub-Gaussian distributions. Furthermore, by construction, the James–Stein (JS) BN yields a smaller local Lipschitz constant compared to the vanilla BN, implying better regularity properties and potentially improved robustness. This facilitates and justifies the application of Stein shrinkage to estimate the mean and variance parameters in BN and the use of it in image classification and segmentation tasks with and without adversarial attacks. We present SOTA performance results using this Stein-corrected BN in a standard ResNet architecture applied to the task of image classification using CIFAR-10 data, 3D CNN on PPMI (neuroimaging) data, and image segmentation using HRNet on Cityscape data with and without adversarial attacks.
Keywords: James-Stein shrinkage estimator, Admissibility, Batch normalization, Sub-Gaussian random variables, Adversarial attacks
@@ 1 Introduction
Deep Neural Networks (DNNs) are now widely accepted and used as the primary workhorse in many of the complicated function approximation tasks encountered in engineering and science. In particular, the field of computer vision has seen a major paradigm shift from traditional innovative feature engineering to data-driven feature learning via the use of DNNs. DNN training is well known to be time intensive and unstable. Batch normalization (BN) is a particularly useful tool in achieving regularization and numerical stability during network training. BN typically involves centering and scaling feature maps. The centering and scaling operations involve computing the sample means (SMs) and variances (SVs) from the samples/members in the batch, respectively. This task is ideally suited for the application of Stein’s shrinkage method, which is well known in the field of statistics. Before delving into Stein’s shrinkage method, let us briefly review the widely popular BN in DNNs.
@@
BN was introduced in the machine learning literature primarily to accelerate training and reduce the so-called internal covariate shift (Ioffe and Szegedy, 2015). Since its inception, several variants have populated the literature, including, but not limited to, layer normalization in Recurrent Neural Networks (RNNs) (Jimmy Lei et al., 2016) and Transformers, spectral normalization in generative adversarial networks (GANs) (Miyato et al., 2018), and many others. Improvements in BN in the form of a stochastic “whitening” transform to decorrelate the features in each layer at each step of the training have also been proposed in Zhang et al. (2021) and BN preconditioning in Susanna et al. (2022), wherein the preconditioner is constructed using batch statistics. These normalization algorithms do not address the effectiveness of theoretical improvements in the stability or convergence of training in the presence of adversarial attacks.
The primary motivation of this paper is to explore whether classical statistical techniques, specifically James–Stein (JS) shrinkage estimators, can bring theoretical and practical benefits when applied to DNNs. Although many advances in adversarial robustness focus on architectural changes or alternative training objectives, our approach is fundamentally different: we revisit a classical estimator from multivariate statistics and investigate how its well-understood risk-reduction properties behave in the presence of adversarial perturbations.
This motivation comes from previous empirical work by Khoshsirat and Kambhamettu (2024), which demonstrated that applying shrinkage estimators to the mean and variance of BN layers improved performance under standard non-adversarial conditions. The improvement can be attributed to the fact that shrinkage estimators reduce the mean squared error (MSE) under squared loss when estimating multiple parameters jointly – a setting that naturally arises in BN, where each layer must estimate a vector of means and variances simultaneously.
In this work, we extend this line of inquiry to adversarial scenarios. Specifically, we address the problem of designing a variant of BN that maintains or improves the prediction accuracy in the presence of adversarial attacks, modeled as additive bounded perturbations (Maurya and Honorio, 2022). Popular attack methods satisfying this bounded perturbation assumption include the fast gradient sign method (FGSM) (Goodfellow et al., 2015) and the iterative variant projected gradient descent (PGD). Unlike many existing defenses that focus on robust losses or architectural modifications, our goal here is not the development of a statistically robust estimator that is insensitive to adversarial attacks by seeking a robust loss or modifying the network architecture, but rather, to adapt an existing shrinkage estimator, specifically, the well known JS-shrinkage estimator and prove that it retains its dominance over the SM and SV in the presence of adversarial perturbations. We theoretically show and empirically confirm that this dominance translates into more stable BN statistics, which in turn are observed to yield improved robustness and prediction accuracy.
1.1 A Brief Note on the JS-shrinkage Estimator
@@
In the following we present a brief introduction to the JS-shrinkage estimator and motivate its aptness in the context of BN in DNNs. This will provide the necessary background and motivation for our work presented in this paper.
[…]
In the case of the JS-estimator, the parameter $\theta$ is a $p$-dimensional mean vector $\boldsymbol{\mu}$ of a Gaussian random vector $\mathbf{X}\sim N(\boldsymbol{\mu},\mathbf{I})$. Under squared error loss, the JS estimator $\boldsymbol{\mu}^{JS}$ dominates the maximum likelihood estimator (MLE) $\boldsymbol{\mu}^{MLE}$, which means that
$\mathbb{E}\!\left[\|\boldsymbol{\mu}^{JS}-\boldsymbol{\mu}\|^{2}\right]\ \leq\ \mathbb{E}\!\left[\|\boldsymbol{\mu}^{MLE}-\boldsymbol{\mu}\|^{2}\right],$
with strict inequality for some $\boldsymbol{\mu}$. Importantly, this improvement concerns the overall estimation risk of the mean vector, not the risk of individual components.
@@
At the foundation of the JS-shrinkage estimator lies Stein’s lemma (also known as Stein’s identity), which characterizes the normal distribution. This lemma is also fundamental to Stein’s method, widely used for distributional comparisons. We will not discuss this well-known method here and instead refer the interested reader to the recent comprehensive survey by Andreas et al. (2023).
@@
Since its inception, the JS shrinkage estimator has been generalized and new optimal shrinkage estimators have been developed for application to many types of distributions such as the location family (Lawrence, 1966), exponential family (Harold M., 1978), gamma distribution (Anirban, 1986), distributions with quadratic variance function (Xianchao and Lawrence, 2016) and work by Fathi et al. (2022) to relax the Gaussian assumption in high dimensions. For a comprehensive survey on Stein’s method and shrinkage estimators, we refer the reader to (Tatsuya, 2024). In addition, more recently, Stein’s shrinkage estimate has been generalized to the multiple mean estimation problem in the context of kernel mean embeddings using convex combinations of empirical mean estimators (Blanchard et al., 2024), and to distributions on manifold-valued random variables (Chun-Hao et al., 2024).
@@ 1.2 Relevance of JS-estimator to BN
Given the above brief introduction to the JS-shrinkage estimator, we are now ready to present it’s relevance to BN in DNNs and how it can be applied to improve the stability and accuracy of predictions. A wide range of normalization techniques have been developed to facilitate the training of DNNs by normalizing intermediate feature distributions and improving the stochastic optimization dynamics. One of the earliest and most influential approaches is BN, introduced in Ioffe and Szegedy (2015), which normalizes intermediate features using statistics computed over a mini-batch, estimating for each channel both a mean and a variance. In the following, we give a brief and precise description of different types of normalization that are in vogue in the literature.
Let $X\in\mathbb{R}^{N\times C\times H\times W}$ be the feature tensor, where $N$ is the batch size, $C$ the number of channels (feature maps), and $H,W$ the spatial dimensions. Many normalization schemes follow the same basic operation:
$\hat{x}_{i}=\frac{x_{i}-\mu_{i}}{\sqrt{\sigma^{2}_{i}+\epsilon}},\quad\mu_{i}=\frac{1}{|S_{i}|}\sum_{k\in S_{i}}x_{k},\quad\sigma_{i}^{2}=\frac{1}{|S_{i}|}\sum_{k\in S_{i}}(x_{k}-\mu_{i})^{2},$
where $S_{i}$ is the subset of tensor elements over which the statistics are computed. The choice of $S_{i}$ defines the normalization type: BN averages over the batch and spatial positions (per channel) (Ioffe and Szegedy, 2015), LayerNorm averages over channels and spatial positions (per sample) (Jimmy Lei et al., 2016), InstanceNorm averages over spatial positions (per sample, per channel) (Ulyanov et al., 2016), and GroupNorm averages over spatial positions and a group of channels (Wu and He, 2018). After normalization, a learnable affine transformation is applied to preserve the representational flexibility: $y_{i}=\gamma_{i}\hat{x_{i}}+\beta_{i},$ where $\gamma_{i},\beta_{i}$ are trainable parameters indexed by coordinates across which we compute the statistics.
@@
Despite the development of alternatives less sensitive to batch size, BN remains widely used due to its strong empirical impact on optimization dynamics (Santurkar et al., 2018) and generalization. A key observation is that BN estimates a large number of mean and variance values simultaneously, one pair for each channel. This can be naturally viewed as a multitask estimation problem: we estimate a vector of means $\mu=(\mu_{1},\dots,\mu_{c})$ and variances $\sigma^{2}=(\sigma^{2}_{1},\dots\sigma^{2}_{c})$, each based on a finite number of observations.
@@
In such settings, JS-shrinkage estimators offer a statistically grounded way to improve estimation by sharing information across tasks, in this case, across channels (feature maps). The classical result suggests that when estimating multiple parameters jointly under squared loss, shrinking individual estimates toward their grand mean can reduce overall error. Applying this idea in the context of BN could lead to more accurate and stable normalization statistics, particularly in small batch or high variance regimes, where the empirical estimates of mean and variance are noisy and benefit from regularization.
@@
Having established adequate motivation for application of Stein’s shrinkage to BN in a DNN, we are now ready to narrate the rest of the story. But before that, a word on what is to follow. Section 2 delves into the presentation of the appropriate shrinkage formulas for the BN. This is followed by the presentation of theoretical results on the dominance of Stein’s shrinkage formulas in the presence of adversarial attacks in Section 3. In Section 4, we present experimental results that compare the performance of the proposed shrinkage to SOTA on several benchmark data sets. Finally, in Section 5, we present the discussion and conclusions.
2 Stein’s Shrinkage Applied to BN
In this section, we apply Stein’s shrinkage to the BN and present formulas for it’s computation in DNNs.
Let $X\sim\mathcal{N}_{p}(\theta,I_{p})$, where $\theta\in\mathbb{R}^{p}$ is an unknown parameter vector and $I_{p}$ is the $p$-dimensional identity matrix. JS-shrinkage estimator of the unknown parameter $\theta$ is given by
$\hat{\theta}^{\mathrm{JS}}=\left(1-\frac{p-2}{\|X\|^{2}_{2}}\right)X.$
@@ 2 Stein’s Shrinkage Applied to BN
It is well known that JS-Estimator has a smaller mean-squared error than the maximum likelihood estimator (MLE) but is obviously a biased estimator. However, a smaller mean squared error is desirable in many applications. Most recently, JS-estimator for the computation of the mean in BN was used in Khoshsirat and Kambhamettu (2024). However, the authors in Khoshsirat and Kambhamettu (2024) used the same JS-estimator formula to shrink the mean as well as the variance in BN. Directly applying the same shrinkage formula to the variance parameter is, however, theoretically flawed, as the SV is not normally distributed but instead follows a scaled chi-squared distribution, a special case of the Gamma distribution. To address this, we adopt a principled alternative.
Much of the theoretical work on BN implicitly relies on the assumption that activation maps that form the input to normalization layers are approximately Gaussian (Lee et al., 2018; Neal, 1994b, a). This Gaussianity assumption enables analytical tractability and underpins many prior analyses of BN dynamics. Following this, we assume that the individual elements of the feature tensor in BN are approximately normally distributed. Under this assumption, the SM is normally distributed, while the SV follows a chi-squared distribution. Specifically, for a given channel with $n=N\cdot H\cdot W$ observations, $\hat{\sigma}_{i}^{2}\sim(\frac{\sigma_{i}^{2}}{n}\chi^{2}_{n-1})\sim\mathrm{Gamma}\left(\alpha=\frac{n-1}{2},\ \beta_{i}=\frac{2\sigma_{i}^{2}}{n}\right)$. Here, $\chi^{2}$ denotes the $\chi$-squared distribution, which is a special case of the well known Gamma distribution with shape parameter $\alpha$ and scale parameter $\beta$. This enables us to apply the shrinkage to the scale parameter of the Gamma distribution using the Stein shrinkage estimator presented in Anirban (1986).
Let $X_{i}\sim\mathrm{Gamma}(\alpha,\beta_{i})$ be independently distributed for $i=1,\dots,p$, where all shape parameters are equal: $\alpha_{1}=\dots=\alpha_{p}=\alpha$. Then, JS-estimators for the scale parameters $\beta=(\beta_{1},\dots,\beta_{p})\in\mathbb{R}_{+}^{p}$ are as follows:
@@
$\hat{\beta}_{i}^{\,\mathrm{JS}}=\frac{X_{i}}{\alpha+1}+\tilde{c}V,\quad\text{where}\quad V=\left(\prod_{j=1}^{p}X_{j}\right)^{1/p},\quad\tilde{c}\in\left[0,\dfrac{2(p-1)}{(\alpha+1)(\alpha p+1)}\right].$ (1)
@@
To incorporate the JS correction into BN, we compute the statistics for each channel and apply the shrinkage formulas as outlined below. Given a 4D tensor input $x\in\mathbb{R}^{N\times C\times H\times W},$ we define $\mu_{C}\in\mathbb{R}^{C}$ and $\sigma^{2}_{C}\in\mathbb{R}^{C}_{+}$ as the channel-wise mean and variance computed over all spatial and batch dimensions:
@@
We then apply shrinkage to these estimates. The JS-corrected mean and variance are:
@@
Equations 2 and 3, respectively, are the formulas we advocate for use in BN corrections and will be used in all our experiments.
We would like to mention that the choice of the shrinkage constant can, in principle, influence the estimator. When the constant approaches zero, the estimator reduces to the empirical variance, so values near the boundary of the admissible interval are of primary theoretical interest. In our experiments, we did not observe a significant sensitivity to moderate changes in this constant. Following the precedent set in the original work on shrinkage estimators for the gamma distribution (Anirban, 1986), we adopted boundary values of the shrinkage constant in our experiments, as detailed in Section 4. We are now ready to present the theoretical results of the dominance of these shrinkage estimators in the presence of adversarial attacks.
3 Dominance of Stein Shrinkage in the Presence of Adversarial Attacks
In this section, we present the main theoretical contribution of our work namely, that the JS-estimator continues to exhibit its dominance even under adversarial perturbations. We begin by formalizing how adversarial attacks can be modeled probabilistically. Then, we introduce the necessary preliminaries and outline the main ideas behind the proof and include the detailed proof in the Appendix A.
We model adversarial perturbations using a sub-Gaussian distribution, which is well motivated both practically and theoretically. Adversarial perturbations are typically bounded in magnitude (for example, under an $l_{p}$-norm constraint) to ensure that the model remains learnable - a standard assumption in the literature (Cohen et al., 2019; Schmidt et al., 2018). Since bounded random variables are inherently sub-Gaussian, this provides a natural modeling choice. Additionally, the sub-Gaussian model enables a tractable analysis via the use of concentration inequalities and allows one to work with a well-defined variance proxy.
Following (Maurya and Honorio, 2022), we assume that adversarial perturbations $Y\in\mathbb{R}^{p}$ are mean-zero sub-Gaussian with variance proxy $2\varepsilon^{2}$, i.e., $\mathbb{E}\!\left[e^{\lambda\langle v,Y\rangle}\right]\leq e^{\varepsilon^{2}\lambda^{2}},\ \forall\lambda\in\mathbb{R},\ \forall v\in\mathbb{R}^{p}:\|v\|=1,$ which we denote compactly as $Y\sim\mathrm{SG}(2\varepsilon^{2})$. This assumption allows us to model the observed vector as a sum of noiseless data and adversarial noise: $Z=X+Y,\quad X\sim\mathcal{N}_{p}(\theta,\sigma^{2}I).$